# DCD: Discriminative and Consistent Representation Distillation

### v3

@@ Abstract
Knowledge Distillation (KD) aims to transfer knowledge from a large teacher model to a smaller student model. While contrastive learning has shown promise in self-supervised learning by creating discriminative representations, its application in knowledge distillation remains limited and focuses primarily on discrimination, neglecting the structural relationships captured by the teacher model. To address this limitation, we propose Discriminative and Consistent Distillation (DCD), which employs a contrastive loss along with a consistency regularization to minimize the discrepancy between the distributions of teacher and student representations. Our method introduces learnable temperature and bias parameters that adapt during training to balance these complementary objectives, replacing the fixed hyperparameters commonly used in contrastive learning approaches. Through extensive experiments on CIFAR-100 and ImageNet ILSVRC-2012, we demonstrate that DCD achieves state-of-the-art performance, with the student model sometimes surpassing the teacher’s accuracy. Furthermore, we show that DCD’s learned representations exhibit superior cross-dataset generalization when transferred to Tiny ImageNet and STL-10. Code is available at https://github.com/giakoumoglou/distillers.
@@ 1 Introduction
Figure 1: Radar plots comparing top-1 accuracy (%) of different distillation methods with DCD on CIFAR-100. Methods are represented by color as follows: Teacher (blue), Student (orange), KD [31] (green), CRD [60] (red), CRD+KD [60] (purple), DCD (ours) (brown), and DCD+KD (ours) (pink).
@@
    We propose DCD, an approach that combines contrastive learning with consistency regularization to ensure both discriminative and structurally consistent representations. Our method eliminates the need for memory banks and introduces learnable parameters that dynamically adjust during training, enabling more efficient knowledge transfer than existing approaches.
@@
    We demonstrate the effectiveness of DCD through extensive experiments on standard benchmarks, showing significant improvements in both accuracy and robustness. DCD outperforms other methods, achieving a 20.31% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{DCD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{DCD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of DCD, KD, and vanilla training of the $i$-th student model, respectively [60]. over the original KD. When combined with KD, it shows a 73.87% relative improvement over the original KD.
The rest of this paper is organized as follows. Section 2 reviews related work in KD and contrastive learning. Section 3 details our proposed DCD methodology. Section 4 presents our experimental setup and results, and Section 5 concludes the paper.
@@ 2 Related Work
Our method bridges the gap between contrastive learning and knowledge distillation by integrating contrastive learning principles with an explicit consistency regularization.
@@ 2.1 Knowledge Distillation
The seminal work on KD [31] introduced the concept of transferring knowledge through softened logit outputs, using temperature scaling in the softmax to better capture the teacher model’s knowledge. Similar to the original KD work, our method utilizes temperature scaling, but we extend this concept through our adaptive temperature mechanism. Rather than using a fixed temperature parameter, DCD learns the optimal temperature during training, allowing for more flexible and dynamic knowledge transfer.
Logit-based methods.
Following this direction, several works have enhanced logit-based distillation through various techniques. Some recent advancements include label decoupling [72], instance-specific label smoothing [66], probability reweighting [49], and normalizing logits before applying softmax and KL divergence [59]. These methods primarily focus on improving how the teacher’s predictions are processed and transferred to guide the student’s learning.
Feature-based methods.
Another line of research focuses on leveraging intermediate feature representations for knowledge transfer. FitNets [55] pioneered this approach by using intermediate ”hints” to guide the student’s learning process. Subsequent works have explored various aspects of feature-based transfer: attention transfer [67] aligns attention maps between teacher and student, correlation congruence [52] preserves structural relationships in feature spaces, and Relational Knowledge Distillation (RKD) [50] transfers mutual relations of data examples. Recent advances include cross-stage connection paths [12], direct reuse of teacher’s classifier [8], and many-to-one representation matching [41]. Notably, Contrastive Representation Distillation (CRD) [60] leverages contrastive learning to maximize mutual information between teacher and student representations, though it requires large memory buffers for storing negative samples. Like RKD, we preserve relational information through our consistency regularization term that explicitly maintains structural relationships. Moreover, our approach eliminates the need for memory banks through efficient in-batch negative sampling and adds structural consistency preservation, making it both more efficient and effective.
Architecture-aware methods.
Recent work has also explored the interaction between network architecture and distillation effectiveness. Methods in this category include differentiable meta-learning for architecture search [19], training-free frameworks for student architecture selection [20], and graph-based methods for architecture adaptation [44]. These approaches aim to optimize not just the distillation process but also the underlying network structures to achieve better knowledge transfer.
@@ 3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Discriminative and Consistent Distillation (DCD), focuses on learning representations that are both discriminative through contrastive learning and structurally consistent with the teacher model through a consistency regularization. DCD ensures that the student model learns to differentiate between different instances while preserving the distributional relationships captured by the teacher model. Figure 2 shows an overview of the proposed method in the latent space.
@@ 3.3 Implementation Details
Memory-efficient sampling.
@@
Adaptive temperature scaling.
@@ 4 Experiments
We evaluate our DCD framework in the KD task for model compression of a large network to a smaller one, similar to [60]. This method aligns with common practices in the field, ensuring a fair comparison between different techniques.
@@ 4.1 Experimental Setup
We implement DCD in PyTorch following the implementation of CRD [60]. The detailed algorithm of DCD is provided in the supplementary material.
Datasets.
@@
We conduct experiments on four popular datasets for model compression: (1) CIFAR-100 [37] contains 50,000 training images with 500 images per class and 10,000 test images. (2) ImageNet ILSVRC-2012 [18] includes 1.2 million images from 1,000 classes for training and 50,000 for validation. (3) STL-10 [17] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (4) Tiny ImageNet [18] comprises 200 classes, each with 500 training images and 50 validation images.
Setup.
@@
We experiment with student-teacher combinations of different capacities, such as ResNet [27] or Wide ResNet (WRN) [68], VGG [58], MobileNet [56], and ShuffleNet [71, 43] (more details are described in the supplementary material). We set $\alpha=0.5$ and $\beta=1$ and ablate both $\alpha$ and $\beta$ in the supplementary material. The hyperparameter $\lambda$ is set to $1.0$ for the KL divergence loss to maintain consistency with [60, 7, 8] (we also ablate $\lambda$ in the supplementary material). Both the student and teacher outputs are projected to a 128-dimensional space using a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. The projection layer, beyond matching dimensions, plays a crucial role in KD by implicitly encoding relational information from previous samples [46]. We empirically set $\tau_{\text{max}}=10.0$. More details on the training procedures can be found in the supplementary material.
Comparison.
@@
We compare our approach to the following state-of-the-art methods: (1) KD [31]; (2) FitNets [55]; (3) AT [67]; (4) SP [61]; (5) CC [52]; (6) VID [1]; (7) RKD [50]; (8) PKT [51]; (9) AB [30]; (10) FT [34]; (11) FSP [65]; (12) NST [33]; (13) CRD [60]; (14) OFD [29]; (15) WSLD [72]; (16) IPWD [49]. In the supplementary material, we include additional methods.
4.2 Results on CIFAR-100
Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing various distillation methods. The values in bold indicate the maximum of each column while underlined values mark the second best. $\Delta$ represents the mean accuracy improvement (%) over the baseline student model.
Teacher WRN-40-2 WRN-40-2 resnet-56 resnet-110 resnet-110 resnet-32x4 VGG-13 $\Delta$
Student WRN-16-2 WRN-40-1 resnet-20 resnet-20 resnet-32 resnet-8x4 VGG-8
Teacher 75.61 75.61 72.34 74.31 74.31 79.42 74.64 -
Student 73.26 71.98 69.06 69.06 71.14 72.50 70.36 0.00
KD [31] 74.92 73.54 70.66 70.67 73.08 73.33 72.98 +1.69
FitNet [55] 73.58 72.24 69.21 68.99 71.06 73.50 71.02 +0.32
AT [67] 74.08 72.77 70.55 70.22 72.31 73.44 71.43 +1.06
SP [61] 73.83 72.43 69.67 70.04 72.69 72.94 72.68 +0.99
CC [52] 73.56 72.21 69.63 69.48 71.48 72.97 70.81 +0.40
VID [1] 74.11 73.30 70.38 70.16 72.61 73.09 71.23 +1.07
RKD [50] 73.35 72.22 69.61 69.25 71.82 71.90 71.48 +0.18
PKT [51] 74.54 73.45 70.34 70.25 72.61 73.64 72.88 +1.48
AB [30] 72.50 72.38 69.47 69.53 70.98 73.17 70.94 +0.20
FT [34] 73.25 71.59 69.84 70.22 72.37 72.86 70.58 +0.45
FSP [65] 72.91 n/a 69.95 70.11 71.89 72.62 70.33 +0.41
NST [33] 73.68 72.24 69.60 69.53 71.96 73.30 71.53 +0.67
CRD [60] 75.48 74.14 71.16 71.46 73.48 75.51 73.94 +2.54
CRD+KD [60] 75.64 74.38 71.63 71.56 73.75 75.46 74.29 +2.48
OFD [29] 75.24 74.33 70.38 n/a 73.23 74.95 73.95 +2.30
WSLD [72] n/a 73.74 71.53 n/a 73.36 74.79 n/a +2.19
IPWD [49] n/a 74.64 71.32 n/a 73.91 76.03 n/a +2.81
DCD (ours) 74.99 73.69 71.18 71.00 73.12 74.23 73.22 +1.72
DCD+KD (ours) 76.06 74.76 71.81 72.03 73.62 75.09 73.95 +2.82
Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods. The values in bold indicate the maximum of each column while underlined values mark the second best. $\Delta$ represents the mean accuracy improvement (%) over the baseline student model.
Teacher VGG-13 ResNet-50 ResNet-50 ResNet-32x4 ResNet-32x4 WRN-40-2 $\Delta$
Student MobileNet-v2 MobileNet-v2 VGG-8 ShuffleNet-v1 ShuffleNet-v2 ShuffleNet-v1
Teacher 74.64 79.34 79.34 79.42 79.42 75.61 -
Student 64.60 64.60 70.36 70.50 71.82 70.50 0.00
KD [31] 67.37 67.35 73.81 74.07 74.45 74.83 +3.25
FitNet [55] 64.14 63.16 70.69 73.59 73.54 73.73 +1.08
AT [67] 59.40 58.58 71.84 71.73 72.73 73.32 -0.80
SP [61] 66.30 68.08 73.34 73.48 74.56 74.52 +2.98
CC [52] 64.86 65.43 70.25 71.14 71.29 71.38 +0.50
VID [1] 65.56 67.57 70.30 73.38 73.40 73.61 +1.91
RKD [50] 64.52 64.43 71.50 72.28 73.21 72.21 +0.96
PKT [51] 67.13 66.52 73.01 74.10 74.69 73.89 +2.83
AB [30] 66.06 67.20 70.65 73.55 74.31 73.34 +2.12
FT [34] 61.78 60.99 70.29 71.75 72.50 72.03 -0.51
NST [33] 58.16 64.96 71.28 74.12 74.68 76.09 +1.15
CRD [60] 69.73 69.11 74.30 75.11 75.65 76.05 +4.43
CRD+KD [60] 69.94 69.54 74.58 75.12 76.05 76.27 +4.69
OFD [29] 69.48 69.04 n/a 75.98 76.82 75.85 +4.83
WSLD [72] n/a 68.79 73.80 75.09 n/a 75.23 +4.24
IPWD [49] n/a 70.25 74.97 76.03 n/a 76.44 +5.43
DCD (ours) 68.35 67.39 73.85 74.26 75.26 74.98 +3.62
DCD+KD (ours) 69.77 70.03 74.08 76.01 76.95 76.51 +5.25
Table 1 and Table 2 provide a comprehensive comparison of top-1 accuracies across various KD methods for both identical and differing architectures between student and teacher models on the CIFAR-100 dataset (extended comparison with more state-of-the-art methods is included in the supplementary material). Specifically, Table 1 focuses on scenarios where the student and teacher share the same architecture, while Table 2 explores settings with differing architectures. Our proposed method, DCD, and its combination with KD, consistently achieve superior performance compared to other distillation objectives, including the original KD. Our method surpasses the teacher network’s performance in both same-architecture (WRN-40-2 to WRN-16-2) and cross-architecture (WRN-40-2 to ShuffleNet-v1) scenarios, achieving accuracy gains of 0.45% and 0.90%, respectively (see 1). As shown in the rightmost column of both tables, DCD, when combined with KD, achieves the highest average improvement over the baseline student model, with +2.82% for same-architecture scenarios and +5.25% for cross-architecture scenarios. The strong performance of both CRD and DCD validates the effectiveness of contrastive learning objectives in knowledge distillation, with our adaptive approach showing consistent improvements over the memory bank-based CRD method.
4.3 Results on ImageNet
Table 3 showcases the top-1 accuracy of student networks that were trained using various distillation methods on ImageNet. The results demonstrate the effectiveness of our method on large-scale datasets, highlighting its ability to distill knowledge from complex models and enhance the performance of student networks. Our approach achieves competitive results, surpassing KD and attaining state-of-the-art performance in ResNet-50 to ResNet-18 distillation. Our approach also shows improvement across different architectures, demonstrating its effectiveness in various distillation scenarios. DCD with KD consistently outperforms the standard KD method across all teacher-student combinations.
Table 3: Test top-1 (%) on ImageNet validation test using various distillation methods. The table compares students and teachers of the same and different architecture. The values in bold indicate the maximum of each column while underlined values mark the second best.
Teacher ResNet-34 ResNet-50 ResNet-50
Student ResNet-18 ResNet-18 MobileNet
Teacher 73.31 76.16 76.16
Student 69.75 69.75 69.63
KD [31] 70.67 71.29 70.49
AT [67] 71.03 71.18 70.18
SP [61] 70.62 71.08 n/a
CC [52] 69.96 n/a n/a
VID [1] n/a 71.11 n/a
RKD [50] 70.40 n/a 68.50
AB [30] n/a n/a 68.89
FT [34] n/a n/a 69.88
FSP [65] 70.58 n/a n/a
NST [33] 70.29 n/a n/a
CRD [60] 71.17 71.25 69.07
OFD [29] 71.03 n/a 71.33
WSLD [72] 72.04 n/a 71.52
IPWD [49] 71.88 n/a 72.65
DCD 71.10 71.38 70.51
DCD+KD 71.71 71.65 71.55
4.4 Capturing Inter-class Correlations
Cross-entropy loss ignores the correlations among class logits in a teacher network, often leading to suboptimal knowledge transfer. By employing ”soft targets”, distillation methods such as those described by [31] have successfully captured these correlations, enhancing student learning. Figure 3 evaluates the effectiveness of various distillation approaches on the CIFAR-100 KD task using WRN-40-2 as the teacher and WRN-40-1 as the student. Specifically, we compare students trained without distillation, with attention transfer [67], with KL divergence [31], and with our proposed DCD method. Our results demonstrate that DCD achieves close correlation alignment between teacher and student logits, as evidenced by the minimal differences in their correlation matrices. However, compared to CRD [60], our method achieves less optimal matching, though it still significantly enhances learning efficiency and lowers error rates. The smaller discrepancies between teacher and student logits clearly indicate that the DCD objective captures a substantial amount of correlation structure in the logit, resulting in lower error rates, although it is surpassed slightly by CRD in achieving the closest match. Moreover, our method also enhances representation learning as it employs a contrastive objective.
(a) Student: vanilla
(b) Student: KD [31]
(c) Student: AT [67]
(d) Student: DCD (ours)
Figure 3: Comparison of correlation matrix differences between teacher and student logits across various distillation methods on the CIFAR-100 task. Subfigures show results for (a) students trained without distillation, (b) with KL divergence [31], (c) with attention transfer (AT) [67], and (d) with our DCD method, highlighting better matching between student’s and teacher’s correlations. Results have been re-implemented according to [60].
4.5 Transferability of Representations
Teacher Student KD AT FitNet CRD CRD+KD DCD DCD+KD

### v4

# Discriminative and Consistent Representation Distillation

@@ Abstract
Knowledge Distillation (KD) aims to transfer knowledge from a large teacher model to a smaller student model. While contrastive learning has shown promise in self-supervised learning by creating discriminative representations, its application in knowledge distillation remains limited and focuses primarily on discrimination, neglecting the structural relationships captured by the teacher model. To address this limitation, we propose Discriminative and Consistent Distillation (DCD), which employs a contrastive loss along with a consistency regularization to minimize the discrepancy between the distributions of teacher and student representations. Our method introduces learnable temperature and bias parameters that adapt during training to balance these complementary objectives, replacing the fixed hyperparameters commonly used in contrastive learning approaches. Through extensive experiments on CIFAR-100 and ImageNet ILSVRC-2012, we demonstrate that DCD achieves state-of-the-art performance, with the student model sometimes surpassing the teacher’s accuracy. Furthermore, we show that DCD’s learned representations exhibit superior cross-dataset generalization when transferred to Tiny ImageNet and STL-10111Code is available at https://github.com/giakoumoglou/distillers..
@@ 1 Introduction
    We propose an approach that combines contrastive learning with consistency regularization to ensure both discriminative and structurally consistent representations. Our method eliminates the need for memory banks and introduces learnable parameters that dynamically adjust during training, enabling more efficient knowledge transfer than existing approaches.
@@
    We demonstrate the effectiveness of our approach through extensive experiments on standard benchmarks, showing significant improvements in both accuracy and robustness. DCD outperforms other methods, achieving a 20.31% relative improvement [^2] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{DCD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{DCD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of DCD, KD, and vanilla training of the $i$-th student model, respectively [66]. over the original KD. When combined with KD, it shows a 73.87% relative improvement over the original KD.
The rest of this paper is organized as follows. Section 2 reviews related work in knowledge distillation and contrastive learning. Section 3 details our proposed methodology. Section 4 presents our experimental setup and results, and Section 5 concludes the paper.
@@ 2 Related Work
Our method combines contrastive learning with knowledge distillation by adding consistency regularization.
@@ 2.1 Knowledge Distillation
The original knowledge distillation work by [33] introduced transferring knowledge through softened logit outputs using temperature scaling in the softmax. Similar to [33], our method uses temperature scaling but learns the optimal temperature during training for better knowledge transfer.
Logit-based Methods
Several methods have improved logit-based distillation through techniques like label decoupling [81], instance-specific label smoothing [73], probability reweighting [55], and normalizing logits before softmax and KL divergence [65]. Some works focus on dynamic temperature adjustment [42] and separate learning of response-based and feature-based distillation [79]. Additional improvements include transformations for better teacher-student alignment [80] and methods for knowledge transfer from stronger teachers [35]. Unlike existing methods such as [42] that use fixed temperature schedules and MLPs, our approach introduces truly learnable temperature parameters that adapt during training, making it more flexible and efficient.
Feature-based Methods
The work by [61] used intermediate feature hints to guide student learning, while [74] aligned attention maps between teacher and student. Several methods have explored structural relationships: [58] preserved feature space relationships, [45] ensured functional consistency, and [25] transferred class-level attention information. The method by [56] transferred mutual relations between data examples. Recent works have introduced cross-stage connection paths [12], direct reuse of teacher’s classifier [8], and many-to-one representation matching [47]. Contrastive Representation Distillation (CRD) [66] used contrastive learning to maximize mutual information between representations but needed large memory buffers. Compared to [66], our approach uses efficient in-batch sampling and adds structural consistency, making it more practical. Our method also differs from previous approaches like [35, 58, 57, 67] in how we define and preserve structural relationships.
Architecture-aware Methods
Recent studies have examined how network architecture affects distillation success. Some works developed meta-learning for architecture search [19], training-free student architecture selection [20], and graph-based architecture adaptation [50]. Methods by [28] and [52] addressed distillation between different architectures through unified feature spaces and intermediate networks respectively. These approaches aim to optimize not just the distillation process but also the underlying network structures to achieve better knowledge transfer.
@@ 3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Discriminative and Consistent Distillation (DCD), focuses on learning representations that are both discriminative through contrastive learning and structurally consistent with the teacher model through a consistency regularization. Our method ensures that the student model learns to differentiate between different instances while preserving the distributional relationships captured by the teacher model. Figure 1 shows an overview of the proposed method in the latent space.
@@ 3.3 Implementation Details
Memory-efficient Sampling
@@
Adaptive Temperature Scaling
@@ 4 Experiments
We evaluate our DCD framework in the KD task for model compression of a large network to a smaller one, similar to [66]. Beyond classification, we also validate our method’s effectiveness on the more challenging object detection task.
@@ 4.1 Experimental Setup
We implement DCD in PyTorch following the implementation of [66].
Datasets
@@
We conduct experiments on four popular datasets for model compression: (1) CIFAR-100 [40] contains 50,000 training images with 500 images per class and 10,000 test images. (2) ImageNet ILSVRC-2012 [18] includes 1.2 million images from 1,000 classes for training and 50,000 for validation. (3) STL-10 [17] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (4) Tiny ImageNet [18] comprises 200 classes, each with 500 training images and 50 validation images. (5) MS-COCO [43] contains 118,000 training images and 5,000 validation images with annotations for object detection.
Setup
@@
We experiment with student-teacher combinations of different capacities, such as ResNet [29] or Wide ResNet (WRN) [75], VGG [64], MobileNet [62], and ShuffleNet [78, 49]. We set $\alpha=0.5$ and $\beta=1$. The hyperparameter $\lambda$ is set to $1.0$ for the KL divergence loss to maintain consistency with [66, 7, 8]. Both the student and teacher outputs are projected to a 128-dimensional space using a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. We empirically set $\tau_{\text{max}}=10.0$.
Comparison
@@
We compare our approach to the following state-of-the-art methods: (1) KD [33]; (2) FitNets [61]; (3) AT [74]; (4) SP [67]; (5) CC [58]; (6) VID [1]; (7) RKD [56]; (8) PKT [57]; (9) AB [32]; (10) FT [37]; (11) FSP [72]; (12) NST [36]; (13) CRD [66]; (14) OFD [31]; (15) WSLD [81]; (16) IPWD [55]; (17) CTKD [42].
Computational Cost
While CRD’s memory bank requires approximately 8MB per class ($16$k features ×$\times 128$ dimensions ×$\times 4$ bytes), DCD’s in-batch sampling needs only 0.13MB total ($256\times 128\times 4$ bytes). This efficiency extends to training time - on a 4-GPU machine, DCD completes ImageNet training in approximately 72 hours compared to CRD’s 88 hours, representing an 18% reduction in training time.
4.2 Main Results
We benchmark our method on image classification and object detection tasks.
Results on CIFAR-100
Table 1 compares top-1 accuracies of various KD methods on CIFAR-100. Our DCD+KD achieves superior performance, surpassing the teacher network by +0.45% in same-architecture (WRN-40-2 to WRN-16-2) and +0.90% in cross-architecture (WRN-40-2 to ShuffleNet-v1) scenarios. The method shows significant improvements over baseline students: +2.82% for same-architecture and +5.25% for cross-architecture pairs, outperforming memory bank-based CRD. DCD alone performs slightly below CRD, which uses a large $16$k-feature memory bank. However, DCD with KD achieves better results by combining complementary objectives: KD’s soft targets provide direct class-level supervision through logit-space KL divergence, while DCD ensures feature-space consistency.
Table 1: Test top-1 accuracy (%) on CIFAR-100 of student networks trained with various distillation methods across different teacher-student architectures. Architecture abbreviations: W: WideResNet, R: ResNet, MN: MobileNet, SN: ShuffleNet. Results adapted from [66]. Results for our method are averaged over five runs.
Same architecture Different architecture
Teacher W-40-2 W-40-2 R-56 R-110 R-110 R-32x4 VGG-13 VGG-13 R-50 R-50 R-32x4 R-32x4 W-40-2
Student W-16-2 W-40-1 R-20 R-20 R-32 R-8x4 VGG-8 MN-v2 MN-v2 VGG-8 SN-v1 SN-v2 SN-v1
Teacher 75.61 75.61 72.34 74.31 74.31 79.42 74.64 74.64 79.34 79.34 79.42 79.42 75.61
Student 73.26 71.98 69.06 69.06 71.14 72.50 70.36 64.60 64.60 70.36 70.50 71.82 70.50
KD [33] 74.92 73.54 70.66 70.67 73.08 73.33 72.98 67.37 67.35 73.81 74.07 74.45 74.83
FitNet [61] 73.58 72.24 69.21 68.99 71.06 73.50 71.02 64.14 63.16 70.69 73.59 73.54 73.73
AT [74] 74.08 72.77 70.55 70.22 72.31 73.44 71.43 59.40 58.58 71.84 71.73 72.73 73.32
SP [67] 73.83 72.43 69.67 70.04 72.69 72.94 72.68 66.30 68.08 73.34 73.48 74.56 74.52
CC [58] 73.56 72.21 69.63 69.48 71.48 72.97 70.81 64.86 65.43 70.25 71.14 71.29 71.38
VID [1] 74.11 73.30 70.38 70.16 72.61 73.09 71.23 65.56 67.57 70.30 73.38 73.40 73.61
RKD [56] 73.35 72.22 69.61 69.25 71.82 71.90 71.48 64.52 64.43 71.50 72.28 73.21 72.21
PKT [57] 74.54 73.45 70.34 70.25 72.61 73.64 72.88 67.13 66.52 73.01 74.10 74.69 73.89
AB [32] 72.50 72.38 69.47 69.53 70.98 73.17 70.94 66.06 67.20 70.65 73.55 74.31 73.34
FT [37] 73.25 71.59 69.84 70.22 72.37 72.86 70.58 61.78 60.99 70.29 71.75 72.50 72.03
FSP [72] 72.91 n/a 69.95 70.11 71.89 72.62 70.33 58.16 64.96 71.28 74.12 74.68 76.09
CRD [66] 75.48 74.14 71.16 71.46 73.48 75.51 73.94 69.73 69.11 74.30 75.11 75.65 76.05
CRD+KD [66] 75.64 74.38 71.63 71.56 73.75 75.46 74.29 69.94 69.54 74.58 75.12 76.05 76.27
OFD [31] 75.24 74.33 70.38 n/a 73.23 74.95 73.95 69.48 69.04 n/a 75.98 76.82 75.8
WSLD [81] n/a 73.74 71.53 n/a 73.36 74.79 n/a n/a 68.79 73.80 75.09 n/a 75.23
IPWD [55] n/a 74.64 71.32 n/a 73.91 76.03 n/a n/a 70.25 74.97 76.03 n/a 76.44
CTKD [42] 75.45 73.93 71.19 70.99 73.52 n/a 73.52 68.46 68.47 n/a 74.78 75.31 75.78
DCD (ours) 74.99 73.69 71.18 71.00 73.12 74.23 73.22 68.35 67.39 73.85 74.26 75.26 74.98
DCD+KD (ours) 76.06 74.76 71.81 72.03 73.62 75.09 73.95 69.77 70.03 74.08 76.01 76.95 76.51
Transferability of Representations
Table 2 compares the top-1 test accuracy of WRN-16-2 (student) distilled from WRN-40-2 (teacher) and evaluated on STL-10 and Tiny ImageNet. The student is trained on CIFAR-100, either directly or via distillation, and serves as a frozen feature extractor with a linear classifier. We assess how well different distillation methods enhance feature transferability. Our results show that DCD, both standalone and combined with KD, significantly improves transferability.
Table 2: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and Tiny ImageNet datasets.
Teacher Student KD [33] AT [74] FitNet [61] CRD [66] CRD+KD [66] DCD DCD+KD